OpenStudy (hi54321):

pls help will fan and medal @slinley ? What is the slope-intercept form of the function that contains the point (3, 4) and has a slope of 2? y =_____ x + ____

OpenStudy (kidrah69):

y=mx+b m=slope

OpenStudy (kidrah69):

here you want to use the point slop equation \[y-y1=m(x-x1)\]

OpenStudy (kidrah69):

we know our slope is 2 and our v-3 and y-4 can you plug them in?

OpenStudy (hi54321):

yes?

OpenStudy (kidrah69):

plug them in show me

OpenStudy (hi54321):

idk how sorry that's why I need help

OpenStudy (kidrah69):

Ok slope=2 plug that in place of M 3 is the X value plug that in place of X1 4 is the Y value plug that in place of Y1

OpenStudy (hi54321):

so y=4 and x=3?

OpenStudy (kidrah69):

\[y-4=2(x-3)\] should look like this

OpenStudy (kidrah69):

now just solve that do you know how?

OpenStudy (hi54321):

no sorry

OpenStudy (hi54321):

@kidrah69 ?

OpenStudy (kidrah69):

we want to get it in y=mx+b form SO lets distribute 2(x-3) \[2\times~x=?\] \[2\times-3=?\]

OpenStudy (hi54321):

-6?

OpenStudy (kidrah69):

right so we have \[y-4=2x-6\] add 4 to both sides

OpenStudy (kidrah69):

-6+4=?

OpenStudy (hi54321):

-2

OpenStudy (kidrah69):

right so y=2x-2

OpenStudy (kidrah69):

that would be the final answer

OpenStudy (hi54321):

can you help with a few more?

OpenStudy (kidrah69):

sure but repost them dont put it on this post